OpenStudy (anonymous):

PLEASEEEE Which of the following is a quadratic function? y = x3 - 3x - 20 y = 3x – x + 4 y = x2 – 4x + 7 y = 2x4 + 9x + 17

OpenStudy (eyust707):

A quadratic function has an x^2 and no high exponent

OpenStudy (eyust707):

we call this the degree of the function. A quadratic function has degree 2.

OpenStudy (eyust707):

y = x3 - 3x - 20 has degree 3 y = 3x – x + 4 has degree 1 y = x2 – 4x + 7 has degree 2 =) y = 2x4 + 9x + 17 has degree 4